OpenStudy (anonymous):

simplify n^2+8n+16/n^2-6n+9 divided by n+4/n-3

OpenStudy (pfenn1):

Is this the correct expression?\[\frac{n^2+8n+16}{n^2-6n+9}\div \frac{n+4}{n-3}\]

OpenStudy (anonymous):

yes.

OpenStudy (pfenn1):

Okay that expression is equivalent to \[\frac{n^2+8n+16}{n^2-6n+9}\ \times\frac{n-3}{n+4}\]

OpenStudy (pfenn1):

Can you factor the expression\[n^2+8n+16=(n+a)(n+b)\]

OpenStudy (anonymous):

i totally dont get how to do any of this. sorry.

OpenStudy (pfenn1):

we know that often a binomial expression like \[n^2+8n+16=(n+a)(n+b)=n^2 +(a+b)n +ab\]so we have to find a and b such that a+b = 8 and ab = 16.

OpenStudy (pfenn1):

Let's try a=4 b=4. \[n^2+8n+16=(n+4)(n+4)\]That works. Now we have to factor the denominator\[n^2-6n+9=(n+c)(n+d) \]where c+d = -6 and cd = 9. Can you do this one?

OpenStudy (anonymous):

So what is the answer?

OpenStudy (pfenn1):

Well, c = -3 and d = -3 works. so \[n^2-6n+9=(n-3)(n-3)=(n-3)^2\] Now we can rewrite the original expression from above as\[\frac{n^2+8n+16}{n^2-6n+9}\ \times\frac{n-3}{n+4}=\frac{(n+4)^2}{(n-3)^2} \times \frac{n-3}{n+4}\]which can be simplified to\[\frac{n+4}{n-3}\]

OpenStudy (anonymous):

thank you:)